OpenStudy (anonymous):

will anyone help me!! pleasee!!! if it takes 4 people to do half a job in 15 days then how many days does it take 10 people to finish the other half of the job? help is very appreciated

OpenStudy (saifoo.khan):

lol, more ppl.. more days? c'mon hashir!

OpenStudy (anonymous):

6

OpenStudy (saifoo.khan):

6 is the correct answer!

OpenStudy (anonymous):

yes i know. thanks method?

OpenStudy (anonymous):

explain please? i need to know how to solve it on my own

OpenStudy (anonymous):

lets go slow. i will wait

OpenStudy (anonymous):

since 4 people finish half a job in 15 days, 4 people finish 1/15 of half a job in 1 day and 1 person finishes 1/60 of half a job in 1 day thus in 1 day 10 people finish 1/6 of half a job and thus in 6 days 10 people finish half a job

OpenStudy (anonymous):

ok i wont wait. there are many methods, but here is what i would do because it will work for other problems. first off 4 people take 15 days to do half the job so it would take them 30 days to do the WHOLE job. (whole job is 1) so each person does \[\frac{1}{30\times 4}=\frac{1}{120}\] of a job per day. now you have ten people. each day they do \[\frac{10}{120}=\frac{1}{12}\] of the job . the whole job is 1. put \[\frac{1}{12}x=\frac{1}{2}\] get \[x=6\] anhhuyalex way is much snappier and i like it better. but for other problems might need to figure out rate of each worker.

OpenStudy (anonymous):

thank you guys!!